How do I solve this word problem:
Two consecutive positive odd numbers have the property that their product plus the square of their difference equals 147. Find the larger of these two numbers.

Answer by ankor@dixie-net.com(22740)

Two consecutive positive odd numbers
x, (x+2)
:
have the property that their product plus the square of their difference equals 147.
x(x+2) + ((x+2) - x)^2 = 147; we know the difference is 2, right?
:
x^2 + 2x + 2^2 = 147
:
x^2 + 2x + 4 - 147 = 0
:
x^2 + 2x - 143 = 0
Factors to:
(x+13)(x-11) = 0
The positive solution
x = 11,
then
the larger number = 13
:
:
Check solution in the statement:
"their product plus the square of their difference equals 147."
(11 * 13) + (13-11)^2 = 147
143 + 2^2 = 147